How much pure antifreeze must be added to 12 L of a 40% solution to obtain a 60% solution. (Remember that pure antifreeze is 100% = 1.) What is the correct question to solve this problem?

Answer by stanbon(75887)

How much pure antifreeze must be added to 12 L of a 40% solution to obtain a 60% solution. (Remember that pure antifreeze is 100% = 1.) What is the correct question to solve this problem?
------
Equation:
active + active = active
0.40*12 + 1.00x = 0.60(12+x)
---
Multiply thru by 100 to get:
40*12 + 100x = 60*12 + 60x
40x = 20*12
x = 6 L (amt. of 100% antifreeze that must be added to the mixture)
--------------------------
Cheers,
Stan H.